Suppose that Australia and Brazil have the outputs per worker in producing sleds and clarinets shown in Table 3. Then Brazil has a: Table 1. Productivity of Brazil and Australia Output per worker Australia Brazil Sleds Clarinets 300 2 200 Select one: O a. Comparative advantage in sleds o b. Absolute advantage in clarinets c. Comparative advantage in clarinets d. Absolute advantage in sleds
Answer #1

Opportunity cost of producing 1 sled (Australia) = 2/300 = 1/150 clarinets

Opportunity cost of producing 1 sled (Brazil) = 1/200 = 1/200 clarinets

Therefore, Brazil has a comparative advantage in sleds (lower opportunity cost)